OpenStudy (anonymous):

For each of the following equtions describe the surface and identify the traces. What does it mean by traces? x^2+y^2-2z=0 Thank you very much for your help!!!

OpenStudy (anonymous):

means the set of all points that fit this equation http://www.wolframalpha.com/input/?i=x^2%2By^2-2z%3D0

OpenStudy (anonymous):

oooh "traces" my fault

OpenStudy (anonymous):

cross section parallel to a plane \(xy\) trace, set \(z=0\) for example

OpenStudy (anonymous):

so in this case the \(xy\) trace is given by the equation \(x^2+y^2=0\) which is a point at the origin of the \(xy\) plane

OpenStudy (anonymous):

and the \(xz\) trace is \(x^2-2z=0\) or \(x^2=2z\) a parabola

OpenStudy (anonymous):

Ok, I understand. Thank you!